E416: A designed experiment has been conducted at three levels (A, B, and C) yielding the following “coded” data:

A B C

6 5 3

3 9 4

5 1 2

2

As a major step in the analysis, one would calculate the degrees of freedom for the “error” sum of squares to be:

A. 7

B. 9

C. 6

D. 3
Answer #1

Solution

levels = k = 3

total observations = N = 9

Error degree of freedom = N-k = 9 - 3 = 6

C. 6